OpenStudy (anonymous):

Find the derivative of the function. y = a7 + cos7 x is this righ? -7cos^6x sinx

OpenStudy (baldymcgee6):

What is a?

hartnn (hartnn):

yes.

hartnn (hartnn):

a must be constant

OpenStudy (anonymous):

a^7+cos^7 X ( X is not with 7)

OpenStudy (baldymcgee6):

Then yes, it's correct :)

hartnn (hartnn):

you are correct.

OpenStudy (baldymcgee6):

you can write that as (cos(x))^7

OpenStudy (baldymcgee6):

for future reference :D

OpenStudy (anonymous):

no its not lol

OpenStudy (anonymous):

im doing an anline quiz

OpenStudy (anonymous):

that wasnt the answer guys

OpenStudy (baldymcgee6):

maybe you didn't type it in right.

hartnn (hartnn):

or using latex, u could write as \[\cos^7x\] `\[\cos^7x\]` what were the options? i don't see any mistake here....

hartnn (hartnn):

u should have put brackets!

hartnn (hartnn):

-7(cos^6 x)( sinx)

OpenStudy (anonymous):

can u guys help me with this question

OpenStudy (baldymcgee6):

How would you do it?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

baldymcgee6

OpenStudy (baldymcgee6):

[e^u]' = e^u * du/dx

OpenStudy (baldymcgee6):

[e^(4tsin(2t))]' = e^(4tsin(2t)) * [4tsin(2t)]'

OpenStudy (anonymous):

okayyy...

OpenStudy (baldymcgee6):

that's not the final answer, you still need to differentiate the rest..

OpenStudy (baldymcgee6):

3pi/4

OpenStudy (anonymous):

whats the derivative of sin2t

OpenStudy (baldymcgee6):

What do you think it is?

OpenStudy (baldymcgee6):

use the chain rule

OpenStudy (anonymous):

-cos2t

OpenStudy (baldymcgee6):

[sin(2t)]' = cos(2t) * [2t]'